OpenStudy (anonymous):

Solve 0.25(7 + 3m) = -0.125 m. Solve 0.25(7 + 3m) = -0.125 m. @Mathematics

OpenStudy (anonymous):

\[m = \frac{-175}{875}\]

OpenStudy (anonymous):

These are the answers to choose from.. -2 2 -40.5 40.5

OpenStudy (anonymous):

Distribute the 0.25 first with 7 and 3m to get 1.75+0.75m=-0.125m --- move variables to the same side.. -0.75 -0.75 --- to get 1.75 = -0.875m --- divide by -0.875 to solve for m to get -2=m